Derive


the MU function from the following TU function:


𝑇𝑈 = 200𝑄 − 25𝑄2 + 𝑄3


From this MU function, draw up a table up to the level of Q where MU becomes negative. i.e (Q = 1, 2, etc). Graph these figures. (10 marks)


1
Expert's answer
2021-04-22T14:45:59-0400
"Solution"

Considering the above "TU" function, marginal utility will be obtained by differentiating "TU" as will be shown below.

"MU=\frac{\delta\ x}{\delta\ y}=200-50q+3q^2"

Below are the calculations obtained from the "MU" function after replacing "1,2,3,4,5,6, and\ 7" in the "MU" function

"1. \ 200 –50 + 3 = 153\\ 
2. \ 200 –100 + 12 = 112 \\
3 .\ 200 –150 + 27 = 77 \\ 
4.\ 200 –200 + 48= 48 \\
5. \ 200 –250 + 75 = 25 \\
6. \ 200 –300 + 108 = 8 \\
7. \ 200 –350 + 147 = –3 \\"

Then the following graph will be obtained from the table above.